OpenStudy (anonymous):

A theatre has 37 seats in the first row, 39 seats in the second row, 41 seats in the third row and so on. A. Can the number of seats in each row be modeled by and arithmetic or geometric sequence B. Write the general term for a sequence a_n that gives the number of seats in row n. C. How many rows are there in row 16?

OpenStudy (anonymous):

its can be write as arithmetic because the difference between 1 row and 2 row is 2

OpenStudy (anonymous):

oh ok how would I write the general term of the sequence \[a_n\] that gives the number of seats

OpenStudy (anonymous):

in row n sorry

OpenStudy (dumbcow):

general form \[A_n = A_1 + d(n-1)\] d = 2 , A1 = 37 \[A_n = 37 + 2(n-1)\]

OpenStudy (anonymous):

\[a_n = 16+2(n-1)\] this would find out how many seats are in row 16

OpenStudy (dumbcow):

what about "n" ?

OpenStudy (dumbcow):

n represents the num of rows ..... 16th row implies n=16 \[A_{16} = 37 + 2(16-1)\]

OpenStudy (anonymous):

wow I feel stupid thanks

OpenStudy (anonymous):

39 time 15 correct would give me how many seats in row 16 right? I'm sorry I'm not good at math at all I just want to understand it.

OpenStudy (anonymous):

I got 67

OpenStudy (dumbcow):

haha ... yes its 67

OpenStudy (anonymous):

yes thank you